A satellite revolves around a planet of radius R in time T. What will be the period of revolution around planet of radius 3R.

Grade:12th pass

1 Answers

Arun
6 years ago
Dear Ravi
 
Use Kepler's law
T² = k * r³
For 1st case
T² = k * R³
Now
For second case
T'² = k * (3R)³
T'² = k* 27 R³
Now
T'² = 27 * T²
T' = 3 \sqrt3 T
 
Regards
Arun (askIITians forum expert)